The eight term of a geometrical progression is –7/32. Find it's common ratio if it's first term is 28.


1
Expert's answer
2021-03-22T10:11:55-0400
"b_8=b_1q^7"

Given "b_1=28, b_8=-\dfrac{7}{32}"


"-\dfrac{7}{32}=28q^7"

"q^7=-\dfrac{1}{128}"

"q=-\dfrac{1}{2}"

The common ratio of the geometrical progression is "-\dfrac{1}{2}."